Subtracting Polynomials:

Perimeter of a rectangle the width of a rectangular play ground is 2x-5 feet, and the length is 3x+9 feet, write a polynomial P(x) that represents the perimeter and then evaluate this perimeter polynomials if x is 4 feet.

Answer by ptaylor(2198)

Width(W) of rectangle=2x-5 ft
Length(L) of rectangle=3x+9 ft
And we know that the Perimeter (P(x)) of rectangle=2W+2L
So our equation is:
2(2x-5)+2(3x+9)=P(x) get rid of parens
4x-10+6x+18=P(x) collect like terms
10x+8=P(x) if x=4, then:
P(4)=10*4+8=48 ft

CK
If x=4, then:
Width(W)=2x-5=2*4-5=3 ft
and Length(L)=3x+9=3*4+9=21
Perimeter (P)=2W+2L=2*3+2*21=6+42=48
48=48

Hope this helps----ptaylor